BODMAS RULE

BODMAS RULE

Grade:8

6 Answers

Vijay Mukati
8 years ago
According to this rule, if you have multiple operating in a single equation like addition, multiplication etc. then the order in which they should be solve will beBrackets, Orders, Division, Multiplication, Addition, Subtraction and this is nothing but the full form of BODMAS.
grenade
8 years ago

BODMAS rule decides the order of operations for add, subtract, multiply, divide, etc as below shown B-Brackets (do all operations contained in the brackets first) O-Orders (powers and square root etc) D-Division M-Multiplication A-Addition S- Subtraction

Let’s see an example and check how BODMASS Rule works 30-(2*6+15/3) +8*3/6

Step1: Brackets 2*6+15/3 = 12+5 = 17

Step2: division 30-17+8*1/2

Step3: multiplication 30-17+4

Step 4: addition and Subtraction 30-17+4=17

20+30-5/4*2+ (1+6) (5-6/4)+9*7 1-6*(2+9)/8

BODMAS Rule is very helpful in solving various algebra math problems.

 

plz aprve

ng29
8 years ago

Easy and simple way to remember BODMAS rule!!

B → Brackets first (parentheses)

O → Of (orders i.e. Powers and Square Roots, Cube Roots, etc.)

DM  Division and Multiplication (start from left to right)

AS → Addition and Subtraction (start from left to right)

 

APPROVE IF USEFUL

Hinal Srivastava
8 years ago
In BODMAS...
B stands for brackets
O for of
DM for division and multiplication
AS for addition and subtraction. 
Nandini Menon
8 years ago
BODMAS is nothing but a small set of rules according to whch you can easily and correctly solve any multi-operational math problem.
B stands for Brackets
O stands for Off  (implying brackets off)
D stands for Division
M stands for Multiplication
A stands for Addition
S stands for Subtraction
Aakash
8 years ago
  1. BODMAS stands for Brackets, Order, Divide, Multiply, Add, Subtract and reminds you in what sequence to carry out the operations of arithmetic. As you can see there are three levels to consider. Any bracketed calculations are carried out with first with any such calculations also following the rules of BODMAS.